A research chemist adds a measured amount of HBr gas to pure water at 25oC and obtains a solution with [H3O+] = 8.3x10-4 M. Calculate [OH-]. Is the solution acidic, basic, or neutral? 


1
Expert's answer
2022-02-28T11:04:52-0500

[H2O] = [H3O+][OH-]

10-14 = 8.3x10-4 M x [OH-]

[OH-] = 10-14 / 8.3x10-4 = 1.2 x 10-11 M

[H3O+] < 10-7 - acidic